2 marbles are drawn without replacement from a box of 3 white 2 green 2 red and 1 blue marble. find the probablity
1. the second marble is blue given the first marble is red. so i got (rd) 1/4 and (blue) 1/7 so 1/4*1/7= 1/28 but my choices are 1/8 b. 1/4 c.1/7 or d. 2/7 so I am not sure what I am doing wrong.
2. the second marble is blue given the first marble is blue and I got 0 since there is only 1 marble is that correct?

Answer by stanbon(75887)

2 marbles are drawn without replacement from a box of 3 white 2 green 2 red and 1 blue marble. find the probablity
1. the second marble is blue given the first marble is red. so i got (rd) 1/4 and (blue) 1/7 so 1/4*1/7= 1/28 but my choices are 1/8 b. 1/4 c.1/7 or d. 2/7 so I am not sure what I am doing wrong.
Ans:P(2nd is blue given the 1st is red) = 2/7
Comment: They don't want the probability of the 2 draw sequence only
the probability getting a blue after a red has been drawn.
----------
2. the second marble is blue given the first marble is blue and I got 0 since there is only 1 marble is that correct?
Ans: CORRECT
===============================
Cheers,
Stan H.